### v1

Abstract
The parallel orbital-updating approach is an orbital iteration based approach for solving eigenvalue problems when many eigenpairs are required, and has been proven to be very efficient, for instance, in electronic structure calculations. In this paper, based on the investigation of a quasi-orthogonality, we present the numerical analysis of the parallel orbital-updating approach for linear eigenvalue problems, including convergence and error estimates of the numerical approximations.
keywords:
parallel orbital-updating, eigenvalue problem, convergence, quasi-orthogonality
{MSCcodes}
65F10, 65J05, 65N25, 65N30
1 Introduction
Eigenvalue problems are typical models in scientific and engineering computing. For instance, Hartree–Fock type equations and Kohn-Sham equations are widely used mathematical models in electronic structure calculations. The eigenvalues and their corresponding eigenfunctions of these equations provide detailed information about the properties of atoms, molecules, and solids, helping to predict chemical reactions, material properties, and physical behaviors (see e.g. [8, 15, 18, 20]).
In electronic structure calculations of a large system, the approximations of a number of eigenpairs are required. With discretization and the self-consistent field iteration [19, 20, 24], solving the Hartree–Fock type equations or the Kohn-Sham equations is then transformed into repeatedly solving some large scale algebraic eigenvalue problems. It is known that the computational cost of solving such large scale eigenvalue problems is huge. In particular, the solving process often requires large scale orthogonalizing operations, which demand global summation operations and limit the large scale parallelization. Nowadays, the computational scale is limited for systems with hundreds to thousands of atoms. Since applications demand and the supercomputers are available, it is significant to develop scalable and parallelizable numerical methods to solve such eigenvalue problems.
To reduce the computational cost and improve the parallel scalability, a so-called parallel orbital-updating (ParO) approach has been proposed in [9] and developed in [11, 21, 22] for solving eigenvalue problems or their equivalent models resulting from electronic structure calculations. With the ParO approach, we avoid solving the large scale eigenvalue problem and instead solve some independent large scale source problems and some small scale projected eigenvalue problems. Moreover, we see from the numerical experiments in [9, 22] that the stiff matrix resulting from the small scale eigenvalue problem is almost diagonal, which may further reduce the computation cost. Because of the independence, these source problems can be solved in parallel intrinsically. For each source problem, the standard parallel strategies can be applied. It then allows a two-level parallelization: one level of parallelization is obtained by partitioning these source problems into different groups of processors, another level of parallelization is obtained by assigning each source problem to several processors contained in each group. This two-level parallelization demonstrates that the ParO approach has a great potential for large-scale calculations. In fact, the numerical experiments in [9, 11, 22] show the effectiveness of the ParO approach. We conclude that the ParO approach is a powerful parallel computing approach to solving eigenvalue problems, in which many eigenpairs are required. However, up to now, there is no any mathematical justification for the ParO approach.
The purpose of this paper is to present the numerical analysis of the ParO approach for linear eigenvalue problems. We see that, orthogonalizing operations in the process of solving eigenvalue problems, for which the computational cost is very expensive, is usually demanded in the scientific and engineering computing such as electronic structure calculations and quite effects on the efficiency and stability of algorithms [20, 29]. We observe that during the implementation process of the ParO approach, we are able to obtain approximately orthogonal orbitals, which we call quasi-orthogonal orbitals. The ParO approach can be viewed as to utilize the quasi-orthogonal approximations, for which the computational cost is less expensive, to obtain orthogonal approximations. Our numerical analysis is starting from the introduction and investigation of a quasi-orthogonality, which plays a crucial role in the orthogonalization of approximations of the eigenvalue problem. We understand that the presence of both single eigenvalues and multiple eigenvalues renders traditional methods for analyzing single eigenvalues no longer applicable. The difficulty for the case of multiple eigenvalues lies in the fact that the traditional measure for the eigenfunction errors is not valid anymore, because the approximate eigenfunctions obtained in iterations may not approximate the same eigenfunction. Instead of focusing on particular eigenfunctions, in our analysis, we employ the eigenspaces and the gap between the eigenspaces, which brings additional analysing complexities and requires sophisticate functional analysis.
Some approaches for constructing source problems in the ParO approach have been proposed in [9]. As a practical example, the shifted-inverse based ParO algorithm applies the shifted-inverse approach to construct some source problems and solves a small scale eigenvalue problem in each iteration to update the shift parameters to speed up the convergence [9, 22]. To analyze the convergence of the algorithm, we first study its simplified version, which fixes the shift parameters and does not carry out the steps of solving small scale eigenvalue problems in iterations. Under the framework of the ParO approach, we show the convergence of numerical solutions produced by the simplified algorithm, which does not require sufficiently accurate initial guesses. Based on the numerical analysis of the simplified version, we then present a more general and informative convergence result of the shifted-inverse based ParO algorithm than the classical results of the shifted-inverse approach for simple eigenvalues mentioned in, e.g., [2, 23]. To improve the numerical stability, a modified version is proposed in [22], which augments the projected subspace by using the residuals. We also provide a brief outline of the proof for the convergence of this modified algorithm.
The rest of this paper is organized as follows. We recall some existing results of a model problem and introduce the relevant notations in Section 2, and provide some elementary analysis for the quasi-orthogonality in Section 3. In Section 4, we carry out numerical analysis for the ParO approach and its several practical versions. Finally, we give some concluding remarks in Section 5.
2 Preliminaries
In this section, we recall some existing results for an eigenvalue problem (including its finite dimensional approximations) that will be used.
[…]
(1) $a(u,v)=\lambda b(u,v),\quad\forall v\in H,$
where $a(\cdot,\cdot)$ and $b(\cdot,\cdot)$ are two symmetric bilinear forms over $H\times H$. We assume that
$a(v,w)\leqslant C_{a}\|v\|\|w\|,\quad\forall v,w\in H,$
and
$a(v,v)\geqslant c_{a}\|v\|^{2},\quad v\in H,$
[…]
It is known that Eq. 1 has a countable sequence of real eigenvalues $0<\lambda_{1}<\lambda_{2}<\cdots$ and $\lambda_{i}$ has the multiplicity $d_{i}(i=1,2,\ldots)$. The indices of $\lambda_{i}$ are $(i,1),\ldots,(i,d_{i})$, that is
$\lambda_{i-1}<\lambda_{i}=\lambda_{i1}=\cdots=\lambda_{id_{i}}<\lambda_{i+1},\quad i=1,2,\ldots,$
with $\lambda_{0}=0,d_{0}=0$.
Define $(i,j)<(r,s)$ if $i<r$ or $i=r,j<s$ with $1\leqslant j\leqslant d_{i}$ and $1\leqslant s\leqslant d_{r}$. Let $M(\lambda_{i})$ denote the eigenspace corresponding to $\lambda_{i}$ and $\{u_{ij}\}_{j=1}^{d_{i}}$ be the orthonormal basis of $M(\lambda_{i})$, that is, $M(\lambda_{i})=\operatorname{span}\{u_{i1},\ldots,u_{id_{i}}\}$ for $i=1,2,\ldots$ with $b(u_{ij},u_{kl})=\delta_{ik}\delta_{jl}$, where $\delta_{ik}$ and $\delta_{jl}$ are the Kronecker delta.
We consider to obtain the smallest $N$ clustered eigenvalues of Eq. 1 and their corresponding eigenfunctions, and assume that there exists $q\in\mathbb{N}_{+}$ such that $\sum_{i=1}^{q}d_{i}=N.$
A typical example of Eq. 1 is an eigenvalue problem of a partial differential operator over a bounded domain. Let $\Omega\subset\mathbb{R}^{d}(d\geqslant 1)$ be a polygonal domain. We shall use the standard notation for Sobolev spaces $H^{1}(\Omega)$ with associated norms (see, e.g. [1]). Let $H=H_{0}^{1}(\Omega)=\{v\in H^{1}(\Omega):v|_{\partial\Omega}=0\}$ and (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) be the standard $L^{2}$ inner product. Consider the eigenvalue problem: find $\lambda\in\mathbb{R}$ and $u\in H_{0}^{1}(\Omega)$ with $\|u\|_{L^{2}(\Omega)}=1$ such that
$-\nabla\cdot(A\nabla u)+cu=\lambda u,$
where $A:\Omega\rightarrow\mathbb{R}^{d\times d}$ is piecewise Lipschitz and symmetric positive definite and $0\leqslant c\in L^{\infty}(\Omega)$. Its associate weak form reads that: find $\lambda\in\mathbb{R}$ and $0\neq u\in H_{0}^{1}(\Omega)$ such that
$a(u,v)=\lambda b(u,v),\quad\forall v\in H_{0}^{1}(\Omega),$
where
$a(u,v)=(A\nabla u,\nabla v)+(cu,v),\quad b(u,v)=(u,v).$
We see that $a(\cdot,\cdot)$ and $b(\cdot,\cdot)$ satisfy the assumptions above.
Remark 2.1.
We mention that the results obtained in this paper are also valid for a more general bilinear form $a(\cdot,\cdot)$ that
$\|v\|^{2}_{H^{1}_{0}(\Omega)}-C_{1}^{-1}\|v\|_{L^{2}(\Omega)}\leqslant C_{2}a(v,v),\quad\forall w\in H^{1}_{0}(\Omega)$
holds for some constant $C_{1},C_{2}>0$ (see, e.g., Remark 2.9 in [12]).
To carry out the analysis, we apply the following distance between two subspaces $U,V\subset H$ ([7, 14, 17]).
$\operatorname{dist}(U,V):=\sup_{u\in U,\|u\|=1}\inf_{v\in V}\|u-v\|.$
Consistently, for any $u,v\in H$, we define
$\operatorname{dist}(u,v):=\operatorname{dist}\left(\operatorname{span}\{u\},\operatorname{span}\{v\}\right).$
It should be noted that $\operatorname{dist}(u,v)$ is actually the sine of the angle between $u$ and $v$, and independent of the norms of vectors.
Note that $\operatorname{dist}(U,V)=1$ when $\operatorname{dim}(U)>\operatorname{dim}(V)$. We also see that for $U,V,W\subset H$ with $\operatorname{dim}(U)=\operatorname{dim}(V)=\operatorname{dim}(W)<\infty$, there holds that
(2) $\operatorname{dist}(U,V)\leqslant\operatorname{dist}(U,W)+\operatorname{dist}(W,V).$
The following useful lemma can be found in [14, 17].
Lemma 2.2.
Given subspaces $U,V\subset H$, if $\operatorname{dim}(U)=\operatorname{dim}(V)<\infty$, then
$\operatorname{dist}(U,V)=\operatorname{dist}(V,U).$
[…]
Let $V^{h}$ be a finite dimensional subspace of $H$ with $\operatorname{dim}(V^{h})=N_{g}$. The standard finite dimensional discretization of Eq. 1 is defined as follows: find $\lambda^{h}\in\mathbb{R}$ and $0\neq u^{h}\in V^{h}$ such that
(3) $a(u^{h},v)=\lambda^{h}b(u^{h},v),\quad\forall v\in V^{h}.$
We may order the eigenvalues of Eq. 3 as follows:
$0<\lambda^{h}_{11}\leqslant\cdots\leqslant\lambda^{h}_{1d_{1}}\leqslant\cdots\leqslant\lambda^{h}_{pd_{p}}.$
We assume $\sum_{i=1}^{p}d_{i}=N_{g}(p\geqslant q,i.e.,N_{g}\geqslant N)$. Indeed, the conclusions in this paper hold for all $N_{g}\geqslant N$. The assumption is adopted to simplify the notations in our numerical analysis. Assume that the corresponding eigenfunctions $u^{h}_{ij}$ for $(i,j)\leqslant(p,d_{p})$ satisfy that $b(u^{h}_{ij},u^{h}_{kl})=\delta_{ik}\delta_{jl}$. For $i=1,\ldots,p$, set $M_{h}(\lambda_{i})=\operatorname{span}\{u^{h}_{i1},\ldots,u^{h}_{id_{i}}\}$.
We obtain from the minimum-maximum principle [4, 7] that
$\lambda_{i}\leqslant\lambda^{h}_{i1}\leqslant\cdots\leqslant\lambda^{h}_{id_{i}},\quad i=1,2,\ldots,p.$
The following conclusion can be found in [13, 16].
Proposition 2.3.
For the eigenvalue problem Eq. 1 and its finite dimensional discretization Eq. 3, there holds that
$0\leqslant\lambda^{h}_{ij}-\lambda_{i}\leqslant\lambda^{h}_{ij}\operatorname{dist}_{a}^{2}(\bigoplus_{i=1}^{q}M(\lambda_{i}),V^{h}),\quad\forall(1,1)\leqslant(i,j)\leqslant(q,d_{q}).$
Given the eigenvalue problem Eq. 1 and its finite dimensional approximation Eq. 3, the following result is classical and can be found in [4, 7, 16].
Proposition 2.4.
If $\operatorname{dist}_{a}\left(\bigoplus_{i=1}^{q}M(\lambda_{i}),V^{h}\right)\ll 1$, then there exists $\hat{u}_{ij}\in M(\lambda_{i})$ such that
$\left\|u^{h}_{ij}-\hat{u}_{ij}\right\|_{a}\leqslant L\operatorname{dist}_{a}(M(\lambda_{i}),V^{h}),\quad\forall(1,1)\leqslant(i,j)\leqslant(q,d_{q}),$
where $L$ is a constant that is independent of $V^{h}$.
Note that Proposition 2.4 tells only that each orthonormal eigenfunction of Eq. 3 approximates some eigenfunction of Eq. 1, which may not be orthonormal each other (see Corollary 2.11 in [10]). However, in practical applications, the approximate property between the orthonormal approximate eigenfunctions and the orthonormal exact eigenfunctions are usually required, which are of structure-preserving and can be used to prevent the accumulation of errors of approximations.
3 Quasi-orthogonality
To carry out the numerical analysis for the ParO approach, we introduce and investigate a quasi-orthogonality, which plays a crucial role in the orthogonalization of approximations of the eigenvalue problem.
Let $\{v_{j}\}_{j=1}^{n}\subset H$ be linearly independent. Consider the Gram-Schmidt orthogonalization of $\{v_{j}\}_{j=1}^{n}$:
  • Choose $\tilde{v}_{1}=v_{1}.$
  • For $j=2,3,\ldots,n,$ set
    (4) $\tilde{v}_{j}=v_{j}-\sum_{l=1}^{j-1}\frac{a(\tilde{v}_{l},v_{j})}{\left\|\tilde{v}_{l}\right\|_{a}^{2}}\tilde{v}_{l}.$
We have the following useful lemma, which tells the properties of the orthogonalization of quasi-orthogonal vectors.
Lemma 3.1.
Given $\theta\in(0,1)$ and $\delta\in\left(0,\frac{(1-\theta)\theta^{n-1}}{2((1+\theta)^{n-1}-\theta^{n-1})}\right)$, let $\{u_{j}\}_{j=1}^{n},\{v_{j}\}_{j=1}^{n}\subset H$ satisfy
(5) $a(u_{i},u_{j})=\delta_{ij},\,\|v_{j}\|_{a}=1,\quad\|u_{j}-v_{j}\|_{a}\leqslant\delta,\quad i,j=1,2,\ldots,n,$
where $\{v_{j}\}^{n}_{j=1}$ is said to be quasi-orthogonal. If $\{\tilde{v}_{j}\}_{j=1}^{n}$ is obtained by the Gram-Schmidt orthogonalization of $\{v_{j}\}_{j=1}^{n}$, then for $j=2,3,\ldots,n$, there holds
$\left\|\tilde{v}_{j}-v_{j}\right\|_{a}\leqslant\frac{1}{\theta}M_{j}\delta,$
where $M_{j}=\frac{2}{\theta^{j-2}}\left((1+\theta)^{j-1}-\theta^{j-1}\right)$.
Proof 3.2.
To obtain the conclusion, we see from the Eq. 4 that it is sufficient to prove
(6) $\sum_{l=1}^{j-1}|a(\tilde{v}_{l},v_{j})|\leqslant M_{j}\delta,\quad\|\tilde{v}_{j}\|_{a}\geqslant\theta,$
for any $j=2,3,\ldots,n$. We prove Eq. 6 by induction.
Note that
$a(v_{1},v_{2})=a(v_{1}-u_{1},v_{2})+a(u_{1},v_{2}-u_{2}),$
which implies
$|a(v_{1},v_{2})|\leqslant\|v_{1}-u_{1}\|_{a}+\|v_{2}-u_{2}\|_{a}\leqslant 2\delta.$
The facts $\tilde{v}_{1}=v_{1},\|v_{1}\|_{a}=\|u_{1}\|_{a}=1$, and
$\tilde{v}_{2}=v_{2}-\frac{a(\tilde{v}_{1},v_{2})}{\|\tilde{v}_{1}\|_{a}^{2}}\tilde{v}_{1}$
yield
$|a(\tilde{v}_{1},v_{2})|\leqslant 2\delta,\quad\|\tilde{v}_{2}\|_{a}\geqslant 1-2\delta,\quad\|\tilde{v}_{2}-v_{2}\|_{a}\leqslant 2\delta.$
A simple calculation shows that $1-2\delta\geqslant\theta$. Hence, Eq. 6 is true when $j=2$.
We assume Eq. 6 is true for $2\leqslant j\leqslant k<n$. We obtain from Eq. 6 that
(7) $\|\tilde{v}_{j}-v_{j}\|_{a}\leqslant\sum_{l=1}^{j-1}\frac{|a(\tilde{v}_{l},v_{j})|}{\|\tilde{v}_{l}\|_{a}}\leqslant\frac{M_{j}\delta}{\theta},\quad j=2,3,\ldots,k.$
Note that Eq. 5 and the identity
$a(\tilde{v}_{l},v_{k+1})=a(\tilde{v}_{l}-v_{l},v_{k+1})+a(v_{l}-u_{l},v_{k+1})+a(u_{l},v_{k+1}-u_{k+1}),\quad l=1,2,\ldots k$
imply
$|a(\tilde{v}_{l},v_{k+1})|$ $\leqslant\|\tilde{v}_{l}-v_{l}\|_{a}+2\delta,\quad l=1,2,\ldots,k.$
Thus we have
$\sum_{l=1}^{k}|a(\tilde{v}_{l},v_{k+1})|\leqslant 2k\delta+\sum_{l=2}^{k}\|\tilde{v}_{l}-v_{l}\|_{a},\quad k\geqslant 2.$
We obtain from Eq. 6 that
$\|\tilde{v}_{k}-v_{k}\|_{a}\leqslant\frac{1}{\theta}\sum_{l=1}^{k-1}|a(\tilde{v}_{l},v_{k})|\leqslant\frac{M_{k}\delta}{\theta}.$
Then we arrive at
$\sum_{l=1}^{k}|a(\tilde{v}_{l},v_{k+1})|\leqslant M_{k+1}\delta,$
where $M_{k+1}=2k+\frac{1}{\theta}\sum_{l=1}^{k}M_{l}$, i.e., $M_{k+1}=\frac{2}{\theta^{k-1}}\left((1+\theta)^{k}-\theta^{k}\right).$
Due to $\tilde{v}_{k+1}=v_{k+1}-\sum_{l=1}^{k}\frac{a(\tilde{v}_{l},v_{k+1})}{\|\tilde{v}_{l}\|_{a}^{2}}\tilde{v}_{l},$ we have
$\|\tilde{v}_{k+1}\|_{a}\geqslant 1-\sum_{l=1}^{k}\frac{|a(\tilde{v}_{l},v_{k+1})|}{\|\tilde{v}_{l}\|_{a}}\geqslant 1-\frac{M_{k+1}\delta}{\theta}.$
Since $M_{k+1}\leqslant\frac{2}{\theta^{n-2}}\left((1+\theta)^{n-1}-\theta^{n-1}\right)$ is true for $2\leqslant k<n$ and
$\delta<\frac{(1-\theta)\theta^{n-1}}{2((1+\theta)^{n-1}-\theta^{n-1})},$
we conclude that $\|\tilde{v}_{k+1}\|_{a}\geqslant\theta,$ which proves Eq. 6.
After a simple calculation, we have the following conclusion.
Corollary 3.3.
Given $\theta\in(0,1)$ and $\delta\in\left(0,\frac{(1-\theta)\theta^{n-1}}{2((1+\theta)^{n-1}-\theta^{n-1})}\right)$, let $\{u_{j}\}_{j=1}^{n},\{v_{j}\}_{j=1}^{n}\subset H$ satisfy
$a(u_{i},u_{j})=\delta_{ij},\|v_{j}\|_{a}=1\quad i,j=1,2,\ldots,n;$
$\|u_{j}-v_{j}\|_{a}\leqslant\delta,\quad~{}j=1,2,\ldots,n.$
Then there exists $\{w_{j}\}_{j=1}^{n}\subset\operatorname{span}\left\{v_{1},\ldots,v_{n}\right\}$ such that
$a(w_{i},w_{j})$ $=\delta_{ij},\quad i,j=1,2,\ldots,n;$
$\operatorname{dist}_{a}\left(u_{j},w_{j}\right)$ $\leqslant\left(1+\frac{2}{\theta^{n-3}}\left(\left(1+\theta\right)^{n-1}-\theta^{n-1}\right)\right)\delta,\quad j=1,2,\ldots,n.$
For given $\varepsilon\in(0,1)$, we consider $U=\operatorname{span}\{u_{1},\ldots,u_{n}\}\subset H$ satisfying $a(u_{i},u_{j})=\delta_{ij}$ and $V\subset H$ satisfying $\operatorname{dim}(V)=\operatorname{dim}(U)$ and $\operatorname{dist}_{a}(U,V)\leqslant\varepsilon$.
Next we show that $\mathcal{P}_{V}|_{U}$ is an isomorphism from $U$ to $V$. Indeed, for $\tilde{u},\hat{u}\in U$ satisfying $\mathcal{P}_{V}\tilde{u}=\mathcal{P}_{V}\hat{u}$, we obtain from
$a(\tilde{u}-\mathcal{P}_{V}\tilde{u},v)=0,\quad v\in V,$
$a(\hat{u}-\mathcal{P}_{V}\hat{u},v)=0,\quad v\in V$
that
$a(\tilde{u}-\hat{u},v)=0,\quad v\in V,$
and $\tilde{u}=\hat{u}$ due to $\operatorname{dist}_{a}(U,V)\leqslant\varepsilon<1$. Hence, $\mathcal{P}_{V}|_{U}$ is an injection and then is isomorphism from $U$ to $V$ since $\operatorname{dim}(V)=\operatorname{dim}(U)$.
Set $v_{j}=\frac{\mathcal{P}_{V}u_{j}}{\left\|\mathcal{P}_{V}u_{j}\right\|_{a}}$ for $j=1,2,\ldots,n$. Since $\mathcal{P}_{V}$ is an isomorphism, we have $V=\operatorname{span}(\{v_{j}\}_{j=1}^{n})$. It shows that
$\mathrel{\phantom{=}}\|u_{j}-v_{j}\|_{a}=\sqrt{\|u_{j}-\mathcal{P}_{V}u_{j}\|_{a}^{2}+\left\|\mathcal{P}_{V}u_{j}-\frac{\mathcal{P}_{V}u_{j}}{\|\mathcal{P}_{V}u_{j}\|_{a}}\right\|_{a}^{2}}$
$\leqslant\sqrt{\operatorname{dist}_{a}^{2}(U,V)+\left(1-\sqrt{1-\operatorname{dist}_{a}^{2}(U,V)}\right)^{2}}\leqslant\sqrt{2-2\sqrt{1-\varepsilon^{2}}}.$
In our analysis, we need to use the inequality $\sqrt{2-2\sqrt{1-\varepsilon^{2}}}\leqslant\frac{(1-\theta)\theta^{n-1}}{2((1+\theta)^{n-1}-\theta^{n-1})},$ which requires that $\varepsilon\in(0,L(\theta,n))$. Here
$L(\theta,n)=\frac{(1-\theta)\theta^{n-1}}{2((1+\theta)^{n-1}-\theta^{n-1})}\sqrt{1-\frac{1}{4}\left(\frac{(1-\theta)\theta^{n-1}}{2((1+\theta)^{n-1}-\theta^{n-1})}\right)^{2}}.$
We see that $L(\theta,n)\in(0,\frac{1}{2})$ since $\frac{(1-\theta)\theta^{n-1}}{2((1+\theta)^{n-1}-\theta^{n-1})}\in(0,\frac{1}{2})$ as $n\geqslant 2$. Consequently, we arrive at the following proposition from Corollary 3.3, which will play a crucial role in our analysis.
Proposition 3.4.
Given $\theta\in(0,1)$ and $\varepsilon\in(0,L(\theta,n))$, if $U,V\subset H$ satisfy
$\operatorname{dim}(U)=\operatorname{dim}(V)=n,\quad\operatorname{dist}_{a}(U,V)\leqslant\varepsilon,$
then for an orthogonal basis $\{u_{j}\}_{j=1}^{n}$ of $U$, there exists $\{w_{j}\}_{j=1}^{n}\subset V$ satisfying
$a(u_{i},u_{j})$ $=a(w_{i},w_{j})=\delta_{ij},\quad i,j=1,2,\ldots,n,$
$\operatorname{dist}_{a}(u_{i},w_{i})$ $\leqslant\left(1+\frac{2}{\theta^{n-3}}\left(\left(1+\theta\right)^{n-1}-\theta^{n-1}\right)\right)\sqrt{2-2\sqrt{1-\varepsilon^{2}}},\quad i=1,2,\ldots,n.$
Consider subspaces $X,Y\subset V^{h}\subset H$ with decompositions as follows:
$V^{h}=\bigoplus_{i=1}^{p}X_{i},\quad X=\bigoplus_{i=1}^{q}X_{i},\quad Y=\sum_{i=1}^{q}Y_{i},$
where $\operatorname{dim}(X)=N$ and $\operatorname{dim}(X_{i})=\operatorname{dim}(Y_{i})=d_{i}$. The following conclusion will be used in our analysis.
Proposition 3.5.
Given $\theta\in(0,1)$, let $\varepsilon\in(0,\min_{1\leqslant i\leqslant q}L(\theta,d_{i}))$ satisfying
(8) $\max_{1\leqslant i\leqslant q}\left(1+\frac{2}{\theta^{d_{i}-3}}\left(\left(1+\theta\right)^{d_{i}-1}-\theta^{d_{i}-1}\right)\right)\sqrt{2-2\sqrt{1-\varepsilon^{2}}}<\frac{1}{\sqrt{N}}.$
If $\max_{i=1,\ldots,q}\operatorname{dist}_{a}(X_{i},Y_{i})<\varepsilon$, then
(9) $Y=\bigoplus_{i=1}^{q}Y_{i}.$
Proof 3.6.
Let $\{x_{ij}\}_{j=1}^{d_{i}}$ be an orthonormal basis of $X_{i}$ with $a(x_{ij},x_{kl})=\delta_{ik}\delta_{jl}$. We obtain from Proposition 3.4 that there exists an orthonormal basis $\{y_{ij}\}_{j=1}^{d_{i}}$ of $Y_{i}(i=1,\ldots,q)$ satisfying that for $j=1,\ldots,d_{i}$
(10) $\operatorname{dist}_{a}(x_{ij},y_{ij})\leqslant\left(1+\frac{2}{\theta^{d_{i}-3}}\left(\left(1+\theta\right)^{d_{i}-1}-\theta^{d_{i}-1}\right)\right)\sqrt{2-2\sqrt{1-\varepsilon^{2}}}\triangleq\tilde{\varepsilon}.$
Set $\{\beta_{rt}^{(ij)}\}$ such that $y_{ij}=\sum_{r=1}^{p}\sum_{t=1}^{d_{r}}\beta_{rt}^{(ij)}x_{rt}$ for $(1,1)\leqslant(i,j)\leqslant(q,d_{q})$ and we have that
$\left(y_{11},\cdots,y_{1d_{1}},\cdots,y_{q1},\cdots,y_{qd_{q}}\right)=\left(x_{11},\cdots,x_{1d_{1}},\cdots,x_{pd_{p}}\right)B_{1},$
where
$B_{1}=\begin{pmatrix}\beta_{11}^{(11)}&\cdots&\beta_{11}^{(1d_{1})}&\cdots&\beta_{11}^{(q1)}&\cdots&\beta_{11}^{(qd_{q})}\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots&\\ \beta_{1d_{1}}^{(11)}&\cdots&\beta_{1d_{1}}^{(1d_{1})}&\cdots&\beta_{1d_{1}}^{(q1)}&\cdots&\beta_{1d_{1}}^{(qd_{q})}\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots&\\ \beta_{p1}^{(11)}&\cdots&\beta_{p1}^{(1d_{1})}&\cdots&\beta_{p1}^{(q1)}&\cdots&\beta_{p1}^{(qd_{q})}\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots&\\ \beta_{pd_{p}}^{(11)}&\cdots&\beta_{pd_{p}}^{(1d_{1})}&\cdots&\beta_{pd_{p}}^{(q1)}&\cdots&\beta_{pd_{p}}^{(qd_{q})}\end{pmatrix}.$
We affirm that matrix
$B_{2}=\begin{pmatrix}\beta_{11}^{(11)}&\cdots&\beta_{11}^{(qd_{q})}\\ \vdots&\ddots&\vdots\\ \beta_{qd_{q}}^{(11)}&\cdots&\beta_{qd_{q}}^{(qd_{q})}\end{pmatrix},$
is strictly diagonally dominant provided $\tilde{\varepsilon}<\frac{1}{\sqrt{N}}$. In fact, we obtain from Eq. 10 that
$\tilde{\varepsilon}\geqslant\operatorname{dist}_{a}(x_{ij},y_{ij})=\operatorname{dist}_{a}(y_{ij},x_{ij})=\frac{\left\|\sum_{r=1}^{p}\sum_{t=1}^{d_{r}}\beta_{rt}^{(ij)}x_{rt}-\beta_{ij}^{(ij)}x_{ij}\right\|_{a}}{\left\|\sum_{r=1}^{p}\sum_{t=1}^{d_{r}}\beta_{rt}^{(ij)}x_{rt}\right\|_{a}}.$
Note that
$\left(\beta_{ij}^{(ij)}\right)^{2}\geqslant\left(1-\tilde{\varepsilon}^{2}\right)\sum_{r=1}^{p}\sum_{t=1}^{d_{r}}\left(\beta_{rt}^{(ij)}\right)^{2},\quad j=1,\ldots,d_{i}$
implies
$\left|\beta_{ij}^{(ij)}\right|$ $\geqslant\sqrt{\left(\frac{1}{\tilde{\varepsilon}^{2}}-1\right)\sum_{(r,t)\neq(i,j)}\left(\beta_{rt}^{(ij)}\right)^{2}}>\sqrt{(N-1)\sum_{(r,t)\neq(i,j)}\left(\beta_{rt}^{(ij)}\right)^{2}}$
$\geqslant\sum_{(r,t)\neq(i,j)}\left|\beta_{rt}^{(ij)}\right|,\quad(1,1)\leqslant(i,j)\leqslant(q,d_{q}).$
It follows from the Gershgorin circle theorem that
$\left|\lambda-\beta_{ij}^{(ij)}\right|\leqslant\sum_{(r,t)\neq(i,j)}\left|\beta_{rt}^{(ij)}\right|,\quad\forall\lambda\in\sigma(B_{2}).$
Consequently, we have
$|\lambda|\geqslant\left|\beta_{ij}^{(ij)}\right|-\sum_{(r,t)\neq(i,j)}\left|\beta_{rt}^{(ij)}\right|>0,\quad(1,1)\leqslant(i,j)\leqslant(q,d_{q}),$
and $\operatorname{rank}(B_{2})=N=\operatorname{rank}(B_{1})$, which completes the proof.
Back to Eq. 3, we turn to estimate the distance between the orthonormal approximate eigenfunctions and the orthonormal exact eigenfunctions.
Theorem 3.7.
If $\operatorname{dist}_{a}\left(\bigoplus_{i=1}^{q}M(\lambda_{i}),V^{h}\right)\ll 1$, then there exists an orthonormal basis $\{u^{o}_{ij}\}$ of $M(\lambda_{i})$ with $b(u^{o}_{ij},u^{o}_{kl})=\delta_{ik}\delta_{jl}$ such that
$\operatorname{dist}_{a}\left(u^{o}_{ij},u^{h}_{ij}\right)\leqslant C\operatorname{dist}_{a}\left(\bigoplus_{i=1}^{q}M(\lambda_{i}),V^{h}\right),\quad(1,1)\leqslant(i,j)\leqslant(q,d_{q}),$
where $C$ is a constant that is independent of $V^{h}$.
Proof 3.8.
For the approximate eigenpairs $\left\{(\lambda^{h}_{ij},u^{h}_{ij})\right\}_{(1,1)\leqslant(i,j)\leqslant(q,d_{q})}$, we obtain from Proposition 2.4 that there exists $\hat{u}_{ij}\in M(\lambda_{i})$ such that
$\left\|u^{h}_{ij}-\hat{u}_{ij}\right\|_{a}\leqslant L\operatorname{dist}_{a}(M(\lambda_{i}),V^{h})\leqslant L\operatorname{dist}_{a}\left(\bigoplus_{i=1}^{q}M(\lambda_{i}),V^{h}\right)\ll 1,\quad(1,1)\leqslant(i,j)\leqslant(q,d_{q}),$
where $L$ is a constant that is independent of $V^{h}$, which yields that
$\left\|\frac{u^{h}_{ij}}{\|u^{h}_{ij}\|_{a}}-\frac{\hat{u}_{ij}}{\|\hat{u}_{ij}\|_{a}}\right\|_{a}=\frac{1}{\sqrt{\lambda^{h}_{ij}}}\left\|u^{h}_{ij}-\hat{u}_{ij}+\left(1-\frac{\|u^{h}_{ij}\|_{a}}{\|\hat{u}_{ij}\|_{a}}\right)\hat{u}_{ij}\right\|_{a}$
$\leqslant$ $\frac{1}{\sqrt{\lambda}_{i}}\left(\left\|u^{h}_{ij}-\hat{u}_{ij}\right\|_{a}+\left|\|\hat{u}_{ij}\|_{a}-\|u^{h}_{ij}\|_{a}\right|\right)\leqslant\frac{2}{\sqrt{\lambda}_{i}}L\operatorname{dist}_{a}\left(\bigoplus_{i=1}^{q}M(\lambda_{i}),V^{h}\right).$
Then it follows from Corollary 3.3 that there exists an orthonormal basis $\{u^{o}_{ij}\}$ of $M(\lambda_{i})$ with $b(u^{o}_{ij},u^{o}_{ik})=\delta_{jk}$ such that
$\operatorname{dist}_{a}\left(u^{o}_{ij},u^{h}_{ij}\right)\leqslant C\operatorname{dist}_{a}\left(\bigoplus_{i=1}^{q}M(\lambda_{i}),V^{h}\right),\quad j=1,\ldots,d_{i},$
where $C$ is a constant that is independent of $V^{h}$. By traversing $i=1,\ldots,q$, the proof is completed.
We see that Theorem 3.7 tells that, for the finite dimensional approximation of an eigenvalue problem, there exists a set of orthogonal eigenfunctions whose distance to the orthogonal approximate eigenfunctions is controlled by the distance of subspaces.
In next section, we will present the approximation between iterative solutions $\left\{\left(\lambda_{ij}^{(n)},u_{ij}^{(n)}\right)\right\}$ produced by the ParO approach and solutions of the discrete problem Eq. 3 with the application of Theorem 3.7. Then we obtain the approximation errors between iterative solutions and solutions of Eq. 1 from Proposition 2.3, Theorem 3.7 and the triangle inequalities
$\left|\lambda_{i}-\lambda_{ij}^{(n)}\right|$ $\leqslant\left|\lambda_{i}-\lambda^{h}_{ij}\right|+\left|\lambda^{h}_{ij}-\lambda_{ij}^{(n)}\right|,$
$\operatorname{dist}_{a}(u_{ij}^{o},u_{ij}^{(n)})$ $\leqslant\operatorname{dist}_{a}(u_{ij}^{o},u^{h,o}_{ij})+\operatorname{dist}_{a}(u^{h,o}_{ij},u_{ij}^{(n)}),$
for $(1,1)\leqslant(i,j)\leqslant(q,d_{q})$, where $\{u^{o}_{ij}\}_{j=1}^{d_{i}}$ and $\{u^{h,o}_{ij}\}_{j=1}^{d_{i}}$ with $b(u^{o}_{ij},u^{o}_{kl})=b(u^{h,o}_{ij},u^{h,o}_{kl})=\delta_{ik}\delta_{jl}$ are orthogonal bases of $M(\lambda_{i})$ and $M_{h}(\lambda_{i})$, respectively.
4 Numerical Analysis
With the quasi-orthogonality, in this section, we carry out the numerical analysis of the ParO approach for clustered eigenvalue problems.
4.1 Algorithm framework
We first recall the framework of the ParO approach for the first $N$ clustered eigenvalues and their corresponding eigenfunctions of Eq. 1, which is stated as Algorithm 1. We mention that Algorithm 1 is indeed a modified version of Algorithm 1.1 in [9].
Algorithm 1 A framework for the ParO approach
  1. 1.
    Given a finite dimensional subspace $V^{h}$ and initial data $\left(\lambda_{k}^{(0)},u_{k}^{(0)}\right)\in\mathbb{R}\times V^{h}$ with $b\left(u_{i}^{(0)},u_{j}^{(0)}\right)=\delta_{ij}(i,j=1,2,\ldots,N)$, let $n=0$.
  2. 2.
    For $k=1,2,\ldots,N$, update each orbital $u_{k}^{(n)}$ in parallel and obtain $u_{k}^{(n+1/2)}$.
  3. 3.
    Construct $U_{n+1}=\operatorname{span}\left\{u_{1}^{(n+1/2)},u_{2}^{(n+1/2)},\ldots,u_{N}^{(n+1/2)}\right\}$.
  4. 4.
    If necessary, find a new basis of $U_{n+1}$ by some procedure and obtain eigenpairs $\left(\lambda_{k}^{(n+1)},u_{k}^{(n+1)}\right)$ by some way, or let $u_{k}^{(n+1)}=u_{k}^{(n+1/2)}$ for $k=1,\ldots,N$.
  5. 5.
    If not converge, let $n=n+1$ and go to $2$.
We see that there are several ways to provide initial data in step 1 of Algorithm 4.1. We may obtain initial data from
  • solving a coarse eigenvalue problem as follows: given a finite dimensional subspace $V^{H}$ of $H$ with $\operatorname{dim}(V^{H})>N$, find $(\lambda^{H},u^{H})\in\mathbb{R}\times V^{H}$ satisfying
    (11) $a\left(u^{H},v\right)=\lambda^{H}b\left(u^{H},v\right)\quad\forall v\in V^{H},$
    to obtain eigenpairs $(\lambda_{k}^{H},u_{k}^{H})$ satisfying $b\left(u_{i}^{H},u_{j}^{H}\right)=\delta_{ij}$ for $i,j=1,2,\ldots,N$ and set $(\lambda_{k}^{(0)},u_{k}^{(0)})=(\lambda_{k}^{H},u_{k}^{H})$ for $k=1,\ldots,N$;
  • neural networks based guesses, which can be obtained from the subspace method based on neural networks [30].
Note that eigenvalue problems are usually resulting from physics. We are able to apply the initial values from physical observation or data. For instance, as mentioned in [9], in electronic structure calculations, we may choose initial data from Gaussian-type orbital, Slater-type orbital and atomic orbital based guesses, and so on.
Since we look for clustered eigenvalues and their corresponding eigenfunctions, we shall consider the approximation of each eigenspace as mentioned in Introduction. We understand that it is not trivial to obtain the multiplicity $d_{i}$ of each eigenvalue $\lambda_{i}$. An effective way to approximate the multiplicities is to cluster the initial guesses $\lambda_{1}^{(0)}\leqslant\lambda_{2}^{(0)}\leqslant\cdots\leqslant\lambda_{N}^{(0)}$. By clustering methods such as Bayesian Information Criterion and Silhouette Method (see e.g., [26, 28]), we can get $q^{\prime}$ clusters with $d_{i}^{\prime}$ eigenpairs in the $i-$th cluster $(i=1,\ldots,q^{\prime})$, that is,
(12) $\left\{\left(\lambda_{ij}^{(0)},u_{ij}^{(0)}\right)\right\}_{i=1,\ldots,q^{\prime},j=1,\ldots,d_{i}^{\prime}}=\left\{\left(\lambda_{k}^{(0)},u_{k}^{(0)}\right)\right\}_{k=1,\ldots,N}.$
In this papaer, we assume that $q^{\prime}=q$ and $d_{i}^{\prime}=d_{i}$ for $i=1,\ldots,q^{\prime}$. Indeed, with a sufficient a priori information of the eigenvalue problem, such an assumption will be likely to hold.
Define $U_{0}=\operatorname{span}\left\{u_{1}^{(0)},u_{2}^{(0)},\ldots,u_{N}^{(0)}\right\}$ and
$\lambda_{ij}^{(n+1)}:=\lambda_{\sum_{r=0}^{i-1}d_{r}+j}^{(n+1)},\quad u_{ij}^{(n+1)}:=u_{\sum_{r=0}^{i-1}d_{r}+j}^{(n+1)},\quad(1,1)\leqslant(i,j)\leqslant(q,d_{q}),\quad n\geqslant 0.$
Set
(13) $U_{n}^{(i)}=\operatorname{span}\left\{u_{i1}^{(n)},\ldots,u_{id_{i}}^{(n)}\right\},\quad i=1,\ldots,q,$
then $U_{n}=\sum_{i=1}^{q}U_{n}^{(i)}$.
To update each orbital in step $2$ and step $3$ of Algorithm 1, as pointed in [9], we can apply the shifted-inverse approach, Chebyshev filtering and so on. Define $\mathcal{F}_{n}:U_{n}\rightarrow U_{n+1}$ as follows
$\mathcal{F}_{n}^{(i)}:=\mathcal{F}_{n}|_{U_{n}^{(i)}}:u_{ij}^{(n)}=u_{\sum_{r=0}^{i-1}d_{r}+j}^{(n)}\mapsto u_{\sum_{r=0}^{i-1}d_{r}+j}^{(n+1/2)}\triangleq u_{ij}^{(n+1/2)},\quad(1,1)\leqslant(i,j)\leqslant(q,d_{q}).$
Set
$U_{n+1/2}^{(i)}:=\{u^{(n+1/2)}_{i1},\dots,u^{(n+1/2)}_{id_{i}}\},\quad i=1,\ldots,q.$
By step $3$, we have $U_{n+1}=\sum_{i=1}^{q}U_{n+1/2}^{(i)}$.
In step $4$ of Algorithm 1, the procedure to update the basis of $U_{n+1}$ can be using the (Gram-Schmidt) orthogonalization or solving a small scale eigenvalue problem as follows: find $(\lambda^{(n+1)},u^{(n+1)})\in\mathbb{R}\times U_{n+1}$ satisfying
(14) $a\left(u^{(n+1)},v\right)=\lambda^{(n+1)}b\left(u^{(n+1)},v\right)\quad\forall v\in U_{n+1},$
to obtain eigenpairs $(\lambda_{ij}^{(n+1)},u_{ij}^{(n+1)})$ with $b\left(u_{ij}^{(n+1)},u_{kl}^{(n+1)}\right)=\delta_{ik}\delta_{jl}$ for $(1,1)\leqslant(i,j),(k.l)\leqslant(q,d_{q})$.
The following theorem shows the approximation errors of eigenpairs when solving a small scale eigenvalue problem is carried out under the assumption that orbitals are approximated well.
Theorem 4.1.
If $\operatorname{dist}_{a}(M_{h}(\lambda_{i}),U_{n_{0}+1/2}^{(i)})\ll 1(i=1,\ldots,q)$, then after solving a small scale eigenvalue problem in $U_{n_{0}+1}=\sum_{i=1}^{q}U_{n_{0}+1/2}^{(i)}$, there exists an orthonormal basis $\{u^{h,o}_{ij}\}_{j=1}^{d_{i}}$ of $M_{h}(\lambda_{i})$ with $b(u^{h,o}_{ij},u^{h,o}_{kl})=\delta_{ik}\delta_{jl}$ such that for $(i,j)\leqslant(q,d_{q})$
(15) $|\lambda_{ij}^{(n_{0}+1)}-\lambda^{h}_{ij}|\leqslant$ $L_{1}\max_{1\leqslant i\leqslant q}\operatorname{dist}_{a}^{2}(M_{h}(\lambda_{i}),U_{n_{0}+1/2}^{(i)}),$
(16) $\operatorname{dist}_{a}(u^{h,o}_{ij},u_{ij}^{(n_{0}+1)})\leqslant$ $L_{2}\max_{1\leqslant i\leqslant q}\operatorname{dist}_{a}(M_{h}(\lambda_{i}),U_{n_{0}+1/2}^{(i)}),$
where $L_{1}$ and $L_{2}$ are constants that are independent of $U_{n_{0}+1/2}$.
Proof 4.2.
We obtain from Proposition 3.5 that
$U_{n_{0}+1}=\bigoplus_{i=1}^{q}U_{n_{0}+1/2}^{(i)}.$
For $\psi\in\bigoplus_{i=1}^{q}M_{h}(\lambda_{i})$ with $\|\psi\|_{a}=1$ and the orthonormal basis $\{v^{h}_{ij}\}_{j=1}^{d_{i}}$ of $M_{h}(\lambda_{i})$ with $a(v^{h}_{ij},v^{h}_{kl})=\delta_{ik}\delta_{jl}$, there exists $\{\alpha_{ij}\}$ satisfying $\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\alpha_{ij}^{2}=1$ such that $\psi=\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\alpha_{ij}v^{h}_{ij}.$ It holds that
(17) $\operatorname{dist}_{a}\left(\psi,U_{n_{0}+1}\right)=\left\|\left(\operatorname{I}-\mathcal{P}_{U_{n_{0}+1}}\right)\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\alpha_{ij}v^{h}_{ij}\right\|_{a}$
$\leqslant$ $\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}|\alpha_{ij}|\left\|\left(\operatorname{I}-\mathcal{P}_{U_{n_{0}+1}}\right)v^{h}_{ij}\right\|_{a}\leqslant\sqrt{\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\left\|\left(\operatorname{I}-\mathcal{P}_{U_{n_{0}+1}}\right)v^{h}_{ij}\right\|_{a}^{2}}$
$=$ $\sqrt{\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\operatorname{dist}_{a}^{2}\left(v^{h}_{ij},U_{n_{0}+1}\right)}\leqslant\sqrt{\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\operatorname{dist}_{a}^{2}\left(M_{h}(\lambda_{i}),U_{n_{0}+1}\right)}$
$\leqslant$ $\sqrt{\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\operatorname{dist}_{a}^{2}\left(M_{h}(\lambda_{i}),U_{n_{0}+1/2}^{(i)}\right)}\leqslant\sqrt{N}\max_{1\leqslant i\leqslant q}\operatorname{dist}_{a}(M_{h}(\lambda_{i}),U_{n_{0}+1/2}^{(i)}),$
and
(18) $\operatorname{dist}_{a}\left(\bigoplus_{i=1}^{q}M_{h}(\lambda_{i}),U_{n_{0}+1}\right)\leqslant\sqrt{N}\max_{1\leqslant i\leqslant q}\operatorname{dist}_{a}(M_{h}(\lambda_{i}),U_{n_{0}+1/2}^{(i)}).$
Moreover, there holds $\lambda^{h}_{ij}<\lambda_{i+1}$ due to $\operatorname{dist}_{a}\left(\bigoplus_{i=1}^{q}M_{h}(\lambda_{i}),U_{n_{0}+1}\right)\ll 1$ for $(1,1)\leqslant(i,j)\leqslant(q,d_{q})$. Hence, the error estimate for Algorithm 1 with solving a small eigenvalue problem in the $n_{0}$-th iteration follows from Eq. 18, Proposition 2.3 and Theorem 3.7 applied to Eq. 14 when $n=n_{0}$.
We will see in the next subsection that the requirements $\operatorname{dist}_{a}(M_{h}(\lambda_{i}),U_{n_{0}+1/2}^{(i)})\ll 1(i=1,\ldots,q)$ can be achieved. We understand that one of the main cost in Algorithm 1 is made by step $4$: the generation and solution of the $N$-dimensional eigenvalue problem. Fortunately, we see from the numerical experiments in [9, 22] that the resulting matrix is almost diagonal: the non-diagonal entries are very small and the computational cost is not very expensive.
4.2 Shifted-inverse based ParO algorithm
In this subsection, we shall study the ParO algorithm for solving the clustered eigenvalue problem when the shifted-inverse approach is applied to update each orbital. The shifted-inverse based ParO algorithm (Algorithm 3), which solves a small scale eigenvalue problem in each iteration to update the shift parameters, has been proposed in [9, 22]. To show the convergence, we first consider a simplified version (Algorithm 2) which fixes the shift parameters and does not carry out the steps of solving projected eigenvalue problems in iterations. Based on the numerical analysis of the simplified version, we then prove that the approximations produced by Algorithm 3 converge rapidly.
We set the shift parameter as any convex combination of $\left\{\lambda_{ij}^{(0)}\right\}_{j=1}^{d_{i}}$ denoted by
$\bar{\lambda}_{i}:=\mathcal{C}_{i}\left(\left\{\lambda_{ij}^{(0)}\right\}_{j=1}^{d_{i}}\right),\quad i=1,\ldots,q.$
For instance, we can choose $\mathcal{C}_{i}\left(\left\{\lambda_{ij}^{(0)}\right\}_{j=1}^{d_{i}}\right)=\frac{1}{d_{i}}\sum_{j=1}^{d_{i}}\lambda_{ij}^{(0)}$.
In our discussion, we assume that the shift parameters are always not equal to the eigenvalues of Eq. 3 in the calculation process. Otherwise, we continue the iterative process on other orbitals while keeping the orbitals unchanged.
Then the shifted-inverse approach $\mathcal{F}_{n}$ writes: for $U_{n}^{(i)}=\operatorname{span}\left\{u_{i1}^{(n)},\ldots,u_{id_{i}}^{(n)}\right\},$ $\mathcal{F}_{n}^{(i)}:=\mathcal{F}_{n}|_{U_{n}^{(i)}}:U_{n}^{(i)}\rightarrow U_{n+1/2}^{(i)}$ with $u_{ij}^{(n+1/2)}=\mathcal{F}_{n}^{(i)}u_{ij}^{(n)}$ for $i=1,\ldots,q,j=1,2,\ldots,d_{i}$ satisfying
$a(u_{ij}^{(n+1/2)},v)-\bar{\lambda}_{i}b(u_{ij}^{(n+1/2)},v)=\bar{\lambda}_{i}b(u_{ij}^{(n)},v),\quad\forall v\in V^{h}.$
The simplified shifted-inverse based ParO algorithm is stated as Algorithm 2. Compared with Algorithm 1, step 4 is no longer carried out in this simplified version.
Algorithm 2 Simplified shifted-inverse based ParO algorithm
  1. 1.
    Given a finite-dimensional space $V^{h}$ and $tol>0$, provide and cluster initial data by Eq. 12, i.e., $\left\{\left(\lambda_{ij}^{(0)},u_{ij}^{(0)}\right)\right\}_{(1,1)\leqslant(i,j)\leqslant(q,d_{q})}\subset\mathbb{R}\times V^{h}$ with $b\left(u_{ij}^{(0)},u_{kl}^{(0)}\right)=\delta_{ik}\delta_{jl}$. Set $\bar{\lambda}_{i}=\mathcal{C}_{i}\left(\left\{\lambda_{ij}^{(0)}\right\}_{j=1}^{d_{i}}\right)$ and let $n=0$.
  2. 2.
    For $(1,1)\leqslant(i,j)\leqslant(q,d_{q})$, find $u_{ij}^{(n+1/2)}\in V^{h}$ in parallel by solving
    (19) $a\left(u_{ij}^{(n+1/2)},v\right)-\bar{\lambda}_{i}b\left(u_{ij}^{(n+1/2)},v\right)=\bar{\lambda}_{i}b\left(u_{ij}^{(n)},v\right)\quad\forall v\in V^{h}.$
  3. 3.
    Set $u_{ij}^{(n+1)}=\frac{u_{ij}^{(n+1/2)}}{\left\|u_{ij}^{(n+1/2)}\right\|_{b}}$. If $\frac{\|u_{ij}^{(n+1)}-u_{ij}^{(n)}\|_{b}}{\|u_{ij}^{(n)}\|_{b}}>tol$, let $n=n+1$ and go to $2$.
If the initial guesses approximate the exact eigenvalues good enough, then the source problems Eq. 19 will be ill-conditioned. We mention that there are approaches to deal with ill-conditioned systems (see e.g., [3, 5, 6, 25, 27]). Indeed, it is quite difficult to solve these ill-conditioned systems well, which will be discussed in our other work. Here we assume that such systems can be well solved.
Algorithm 2 may be viewed as an extension of the shifted-inverse approach to clustered eigenvalue problems.
Proposition 4.3.
If $U^{(i)}_{n+1/2}$ is obtained by Algorithm 2, then
$\operatorname{dim}\left(U_{n+1/2}^{(i)}\right)=\operatorname{dim}\left(U_{n}^{(i)}\right),\quad i=1,2,\ldots,q.$
Proof 4.4.
For the $n$-th iteration, consider the linear operators
$\mathcal{F}_{n}^{(i)}:U_{n}^{(i)}\rightarrow U_{n+1/2}^{(i)},\quad i=1,2,\ldots,q,$
and for $u^{(n)}\in U_{n}^{(i)}$, $u^{(n+1/2)}=\mathcal{F}_{n}^{(i)}u^{(n)}$ satisfying
$a\left(u^{(n+1/2)},v\right)-\bar{\lambda}_{i}b\left(u^{(n+1/2)},v\right)=\bar{\lambda}_{i}b\left(u^{(n)},v\right),\quad\forall v\in V^{h}.$
We claim that $\mathcal{F}_{n}^{(i)}$ is an injection. Indeed, $u^{(n+1/2)}=v^{(n+1/2)}$ implies that
$\bar{\lambda}_{i}b\left(u^{(n)},v\right)$ $=a\left(u^{(n+1/2)},v\right)-\bar{\lambda}_{i}b\left(u^{(n+1/2)},v\right)$
$=a\left(v^{(n+1/2)},v\right)-\bar{\lambda}_{i}b\left(v^{(n+1/2)},v\right)=\bar{\lambda}_{i}b\left(v^{(n)},v\right),\quad v\in V^{h},$
and $u^{(n)}=v^{(n)}$.
Since $U_{n}^{(i)}$ and $U_{n+1/2}^{(i)}$ are finite dimensional, $\mathcal{F}_{n}^{(i)}$ is indeed an isomorphism and we arrive at
$\operatorname{dim}\left(U_{n+1/2}^{(i)}\right)=\operatorname{dim}\left(U_{n}^{(i)}\right),\quad i=1,2,\ldots,q.$
With the help of Proposition 4.3, we obtain convergence of eigenspace approximations produced by Algorithm 2.
Theorem 4.5.
Assume that
(20) $0<\delta_{0}:=\max_{(1,1)\leqslant(i,j)\leqslant(q,d_{q})}|\lambda^{h}_{ij}-\bar{\lambda}_{i}|<\frac{g}{2},$
where $g:=\min_{1\leqslant i<r\leqslant q+1}|\lambda^{h}_{id_{i}}-\lambda^{h}_{r1}|;$
(21) $\operatorname{dim}\left(U_{0}^{(i)}\right)=d_{i},\quad\operatorname{dist}_{a}\left(M_{h}(\lambda_{i}),U_{0}^{(i)}\right)<1\quad\forall i=1,2,\ldots,q.$
If $U_{n+1/2}^{(i)}$ is produced by Algorithm 2, then
$\operatorname{dist}_{a}\left(M_{h}(\lambda_{i}),U_{n+1/2}^{(i)}\right)\leqslant\varepsilon_{n+1}=\frac{\delta_{0}\varepsilon_{n}}{\sqrt{(g-\delta_{0})^{2}(1-\varepsilon_{n}^{2})+\delta_{0}^{2}\varepsilon^{2}_{n}}},\quad\lim_{n\rightarrow\infty}\frac{\varepsilon_{n+1}}{\varepsilon_{n}}=\frac{\delta_{0}}{g-\delta_{0}}.$
Proof 4.6.
Let us consider $n=0$ first.
Since $\{u^{h}_{ij}\}_{j=1}^{d_{i}}$ is the orthonormal basis of $M_{h}(\lambda_{i})(i=1,2,\ldots,p)$ with $b(u^{h}_{ij},u^{h}_{kl})=\delta_{ik}\delta_{jl}$, for $v_{i}^{(0)}\in U_{0}^{(i)}$, there exists $\left\{\alpha_{rt}^{(i)}\right\}$ such that
(22) $v_{i}^{(0)}=\sum_{r=1}^{p}\sum_{t=1}^{d_{r}}\alpha_{rt}^{(i)}u^{h}_{rt},\quad i=1,2,\ldots,q.$
A simple calculation and
$a(\mathcal{P}_{M_{h}(\lambda_{i})}v_{i}^{(0)},v)=a(v_{i}^{(0)},v),\quad v\in M_{h}(\lambda_{i}),$
show that
$\mathcal{P}_{M_{h}(\lambda_{i})}v_{i}^{(0)}=\sum_{t=1}^{d_{i}}\alpha_{it}^{(i)}u^{h}_{it}.$
We obtain from Lemma 2.2 and Eq. 21 that there exists $\varepsilon_{0}\in(0,1)$ such that
$\varepsilon_{0}\geqslant$ $\operatorname{dist}_{a}\left(U_{0}^{(i)},M_{h}(\lambda_{i})\right)\geqslant\operatorname{dist}_{a}\left(v_{i}^{(0)},M_{h}(\lambda_{i})\right)$
$=$ $\operatorname{dist}_{a}\left(v_{i}^{(0)},\mathcal{P}_{M_{h}(\lambda_{i})}v_{i}^{(0)}\right)=\frac{\left\|\sum_{r=1}^{p}\sum_{t=1}^{d_{r}}\alpha_{rt}^{(i)}u^{h}_{rt}-\sum_{t=1}^{d_{i}}\alpha_{it}^{(i)}u^{h}_{it}\right\|_{a}}{\left\|\sum_{r=1}^{p}\sum_{t=1}^{d_{r}}\alpha_{rt}^{(i)}u^{h}_{rt}\right\|_{a}},$
which yields,
(23) $\sum_{t=1}^{d_{i}}\lambda^{h}_{it}\left(\alpha_{it}^{(i)}\right)^{2}\geqslant\left(\frac{1-\varepsilon^{2}_{0}}{\varepsilon_{0}^{2}}\right)\sum_{1\leqslant r\neq i\leqslant p}\sum_{t=1}^{d_{r}}\lambda^{h}_{rt}\left(\alpha_{rt}^{(i)}\right)^{2},\quad i=1,2,\ldots,q.$
Let $v_{i}^{(1/2)}\in V^{h}$ satisfy
$a\left(v_{i}^{(1/2)},v\right)-\bar{\lambda}_{i}b\left(v_{i}^{(1/2)},v\right)=\bar{\lambda}_{i}b\left(v_{i}^{(0)},v\right)\quad\forall v\in V^{h}.$
We may write $v_{i}^{(1/2)}=\sum_{r=1}^{p}\sum_{t=1}^{d_{r}}\beta_{rt}^{(i)}u^{h}_{rt}$ for $i=1,2,\ldots,q$. Note that Eq. 3, Eq. 19 and Eq. 22 imply
$\bar{\lambda}_{i}\sum_{r=1}^{p}\sum_{t=1}^{d_{r}}\alpha_{rt}^{(i)}b(u^{h}_{rt},v)=\bar{\lambda}_{i}b(v_{i}^{(0)},v)=a\left(v_{i}^{(1/2)},v\right)-\bar{\lambda}_{i}b\left(v_{i}^{(1/2)},v\right)$
$=$ $\sum_{r=1}^{p}\sum_{t=1}^{d_{r}}\left(\lambda^{h}_{rt}-\bar{\lambda}_{i}\right)\beta_{rt}^{(i)}b(u^{h}_{rt},v),\quad v\in V^{h}.$
We have
$\beta_{rt}^{(i)}=\frac{\bar{\lambda}_{i}}{\lambda^{h}_{rt}-\bar{\lambda}_{i}}\alpha_{rt}^{(i)},\quad i=1,2,\ldots,q,$
and hence,
$v_{i}^{(1/2)}=\sum_{r=1}^{p}\sum_{t=1}^{d_{r}}\frac{\bar{\lambda}_{i}}{\lambda^{h}_{rt}-\bar{\lambda}_{i}}\alpha_{rt}^{(i)}u^{h}_{rt},\quad i=1,2,\ldots,q.$
Note that Eq. 20 and Eq. 23 imply that
$\left|\lambda^{h}_{rt}-\bar{\lambda}_{i}\right|\geqslant$ $\left|\lambda^{h}_{rt}-\lambda^{h}_{i1}\right|-\left|\lambda^{h}_{i1}-\bar{\lambda}_{i}\right|\geqslant g-\delta_{0},\quad r\neq i,$
and
$\frac{\sum_{1\leqslant r\neq i\leqslant p}\sum_{t=1}^{d_{r}}\left(\frac{\bar{\lambda}_{i}}{\lambda^{h}_{rt}-\bar{\lambda}_{i}}\alpha_{rt}^{(i)}\right)^{2}\lambda^{h}_{rt}}{\sum_{t=1}^{d_{i}}\left(\frac{\bar{\lambda}_{i}}{\lambda^{h}_{it}-\bar{\lambda}_{i}}\alpha_{it}^{(i)}\right)^{2}\lambda^{h}_{it}}$
$\leqslant$ $\left(\frac{\delta_{0}}{g-\delta_{0}}\right)^{2}\frac{\sum_{1\leqslant r\neq i\leqslant p}\sum_{t=1}^{d_{r}}\lambda^{h}_{rt}\left(\alpha_{rt}^{(i)}\right)^{2}}{\sum_{t=1}^{d_{i}}\lambda^{h}_{it}\left(\alpha_{it}^{(i)}\right)^{2}}\leqslant\left(\frac{\delta_{0}}{g-\delta_{0}}\right)^{2}\frac{\varepsilon_{0}^{2}}{1-\varepsilon_{0}^{2}}.$
We then get that
(24) $\operatorname{dist}_{a}\left(v_{i}^{(1/2)},M_{h}(\lambda_{i})\right)=$ $\frac{\left\|v_{i}^{(1/2)}-\mathcal{P}_{M_{h}(\lambda_{i})}v_{i}^{(1/2)}\right\|_{a}}{\left\|v_{i}^{(1/2)}\right\|_{a}}$
$=$ $\sqrt{1-\frac{\sum_{t=1}^{d_{i}}\left(\frac{\bar{\lambda}_{i}}{\lambda^{h}_{it}-\bar{\lambda}_{i}}\alpha_{it}^{(i)}\right)^{2}\lambda^{h}_{it}}{\sum_{r=1}^{p}\sum_{t=1}^{d_{r}}\left(\frac{\bar{\lambda}_{i}}{\lambda^{h}_{rt}-\bar{\lambda}_{i}}\alpha_{rt}^{(i)}\right)^{2}\lambda^{h}_{rt}}}$
$\leqslant$ $\frac{\delta_{0}\varepsilon_{0}}{\sqrt{(g-\delta_{0})^{2}(1-\varepsilon_{0}^{2})+\delta_{0}^{2}\varepsilon^{2}_{0}}}\triangleq\varepsilon_{1},\quad i=1,\ldots,q.$
We see that $v_{i}^{(1)}=\mathcal{F}_{0}^{(i)}v_{i}^{(0)}$, where $\mathcal{F}_{0}^{(i)}$ is an isomorphism. Then we obtain from Eq. 20 and Eq. 24 that $\varepsilon_{1}\leqslant\varepsilon_{0}<1$ and
$\operatorname{dist}_{a}\left(M_{h}(\lambda_{i}),U_{1/2}^{(i)}\right)\leqslant\varepsilon_{1},\quad i=1,2,\ldots,q.$
Similarly, we have
$\operatorname{dist}_{a}\left(M_{h}(\lambda_{i}),U_{n+1/2}^{(i)}\right)\leqslant\varepsilon_{n+1}=\frac{\delta_{0}\varepsilon_{n}}{\sqrt{(g-\delta_{0})^{2}(1-\varepsilon_{n}^{2})+\delta_{0}^{2}\varepsilon^{2}_{n}}},\quad\forall n\in\mathbb{N}.$
Thus, $\varepsilon_{n+1}\leqslant\varepsilon_{n},\forall n\in\mathbb{N}$ and
$\varepsilon_{n+1}=$ $\frac{\delta_{0}\varepsilon_{n}}{\sqrt{(g-\delta_{0})^{2}(1-\varepsilon_{n}^{2})+\delta_{0}^{2}\varepsilon^{2}_{n}}}\leqslant\left(\frac{\delta_{0}}{\sqrt{(g-\delta_{0})^{2}(1-\varepsilon_{n-1}^{2})+\delta_{0}^{2}\varepsilon^{2}_{n-1}}}\right)^{2}\varepsilon_{n-1}$
$\leqslant$ $\cdots\leqslant\left(\frac{\delta_{0}}{\sqrt{(g-\delta_{0})^{2}(1-\varepsilon_{0}^{2})+\delta_{0}^{2}\varepsilon^{2}_{0}}}\right)^{n}\varepsilon_{0},$
which indicates that $\varepsilon_{n+1}$ decreases towards 00 as $n\rightarrow\infty$. Moreover, there holds that
(25) $\lim_{n\rightarrow\infty}\frac{\varepsilon_{n+1}}{\varepsilon_{n}}=\frac{\delta_{0}}{g-\delta_{0}}.$
Theorem 4.5 shows that convergence of Algorithm 2 does not require sufficiently accurate initial guesses. Eq. 20 ensures that the shift parameter is closer to the eigenvalue being approximated. Indeed, Eq. 20 can be satisfied under the assumption that the finite dimensional discretization Eq. 3 approximates Eq. 1 not “too badly”. Eq. 21 guarantees that the dimension of the approximated subspace is preserved. When $N=1$, Eq. 25 may be reviewed as the classical shift-inverse convergence result.
We now analyze the convergence of the shifted-inverse based ParO algorithm proposed in [9, 22], which is stated as Algorithm 3.
Algorithm 3 Shifted-inverse based ParO algorithm
  1. 1.
    Given a finite-dimensional space $V^{h}$ and $tol>0$, provide and cluster initial data by Eq. 12, i.e., $\left\{\left(\lambda_{ij}^{(0)},u_{ij}^{(0)}\right)\right\}_{(1,1)\leqslant(i,j)\leqslant(q,d_{q})}\subset\mathbb{R}\times V^{h}$ with $b\left(u_{ij}^{(0)},u_{kl}^{(0)}\right)=\delta_{ik}\delta_{jl}$. Set $\bar{\lambda}_{i}^{(0)}=\mathcal{C}_{i}\left(\left\{\lambda_{ij}^{(0)}\right\}_{j=1}^{d_{i}}\right)$ and let $n=0$.
  2. 2.
    For $(1,1)\leqslant(i,j)\leqslant(q,d_{q})$, find $u_{ij}^{(n+1/2)}\in V^{h}$ in parallel by solving
    (26) $a\left(u_{ij}^{(n+1/2)},v\right)-\bar{\lambda}_{i}^{(n)}b\left(u_{ij}^{(n+1/2)},v\right)=\bar{\lambda}_{i}^{(n)}b\left(u_{ij}^{(n)},v\right)\quad\forall v\in V^{h}.$
  3. 3.
    Construct $U_{n+1}=\operatorname{span}\left\{u_{11}^{(n+1/2)},\ldots,u_{1d_{1}}^{(n+1/2)},\ldots,u_{qd_{q}}^{(n+1/2)}\right\}$.
  4. 4.
    Solve an eigenvalue problem: find $(\lambda^{(n+1)},u^{(n+1)})\in\mathbb{R}\times U_{n+1}$ satisfying
    (27) $a\left(u^{(n+1)},v\right)=\lambda^{(n+1)}b\left(u^{(n+1)},v\right)\quad\forall v\in U_{n+1},$
    to obtain eigenpairs $\left\{\left(\lambda_{ij}^{(n+1)},u_{ij}^{(n+1)}\right)\right\}$ with $b\left(u_{ij}^{(n+1)},u_{kl}^{(n+1)}\right)=\delta_{ik}\delta_{jl}$.
  5. 5.
    If $\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\left|\lambda_{ij}^{(n+1)}-\lambda_{ij}^{(n)}\right|>tol$, set $\bar{\lambda}_{i}^{(n+1)}=\mathcal{C}_{i}\left(\left\{\lambda_{ij}^{(n+1)}\right\}_{j=1}^{d_{i}}\right)$, $n=n+1$ and go to $2$.
As mentioned above, we assume that the shift parameters are always not equal to the eigenvalues of Eq. 3 in the calculation process. If there are cases where some orbitals are very well approximated, while other orbitals have not yet converged, then we continue the iterative process on the non-converged orbitals while keeping the well approximated orbitals unchanged.
The following theorem tells the convergence of Algorithm 3.
Theorem 4.7.
Assume that there exists $0<\varepsilon_{0}\ll 1$ and an orthonormal basis $\{u^{h,o,0}_{ij}\}_{j=1}^{d_{i}}$ of $M_{h}(\lambda_{i})(i=1,\ldots,q)$ with $b(u^{h,o,0}_{ij},u^{h,o,0}_{kl})=\delta_{ik}\delta_{jl}$ such that
(28) $\operatorname{dist}_{a}\left(u^{h,o,0}_{ij},u_{ij}^{(0)}\right)$ $\leqslant\varepsilon_{0},\quad(1,1)\leqslant(i,j)\leqslant(q,d_{q}),$
(29) $\zeta_{0}:=\max_{1\leqslant i\leqslant q}\left|\mathcal{C}_{i}\left(\left\{\lambda^{h}_{ij}\right\}_{j=1}^{d_{i}}\right)-\bar{\lambda}_{i}^{(0)}\right|$ $\ll g,\quad\gamma:=\max_{1\leqslant i\leqslant q}\left(\lambda^{h}_{id_{i}}-\lambda^{h}_{i1}\right)\ll g.$
If $\{u_{ij}^{(n+1)}\}$ are produced by Algorithm 3, then there exists an orthonormal basis $\{u^{h,o,n+1}_{ij}\}_{j=1}^{d_{i}}$ of $M_{h}(\lambda_{i})(i=1,\ldots,q)$ with $b(u^{h,o,n+1}_{ij},u^{h,o,n+1}_{kl})=\delta_{ik}\delta_{jl}$ such that
$\operatorname{dist}_{a}\left(u^{h,o,n+1}_{ij},u_{ij}^{(n+1)}\right)$ $\leqslant\varepsilon_{n+1}=\frac{C\sqrt{DN}\left(\gamma+\zeta_{n}\right)\varepsilon_{n}}{\sqrt{(g-\gamma-\zeta_{n})^{2}(1-D\varepsilon_{n}^{2})+D\left(\gamma+\zeta_{n}\right)^{2}\varepsilon^{2}_{n}}},$
$|\lambda^{h}_{ij}-\lambda_{ij}^{(n+1)}|$ $\leqslant\zeta_{n+1}:=\frac{\lambda^{h}_{q+1,1}}{C^{2}}\varepsilon_{n+1}^{2},\quad(1,1)\leqslant(i,j)\leqslant(q,d_{q}),$
where $D:=\max_{1\leqslant i\leqslant q}d_{i}$ and $C$ is a constant that comes from the application of Theorem 3.7, and is independent of $U_{n}(n=0,1,2,\ldots)$, and
(30) $\lim_{n\rightarrow\infty}\frac{\varepsilon_{n+1}}{\varepsilon_{n}}=\frac{C\sqrt{DN}\gamma}{g-\gamma}.$
Proof 4.8.
Let us start by $n=0$.
Since $\{u^{h,o,0}_{ij}\}_{i=1}^{d_{i}}$ is an orthonormal basis of $M_{h}(\lambda_{i})$, for $\varphi\in M_{h}(\lambda_{i})$ with $\|\varphi\|_{a}=1$, there exists $\{\alpha_{j}\}$ satisfying $\varphi=\sum_{j=1}^{d_{i}}\alpha_{j}u^{h,o,0}_{ij}$ and $\sum_{j=1}^{d_{i}}\alpha_{j}^{2}=1$. Note that
$\operatorname{dist}_{a}\left(\varphi,U_{0}^{(i)}\right)=\left\|\left(\operatorname{I}-\mathcal{P}_{U_{0}^{(i)}}\right)\varphi\right\|_{a}\leqslant\sum_{j=1}^{d_{i}}|\alpha_{j}|\left\|\left(\operatorname{I}-\mathcal{P}_{U_{0}^{(i)}}\right)u^{h,o,0}_{ij}\right\|_{a}$
$=$ $\sum_{j=1}^{d_{i}}|\alpha_{j}|\operatorname{dist}_{a}\left(u^{h,o,0}_{ij},U_{0}^{(i)}\right)\leqslant\sum_{j=1}^{d_{i}}|\alpha_{j}|\operatorname{dist}_{a}\left(u^{h,o,0}_{ij},u_{ij}^{(0)}\right)\leqslant\sqrt{d_{i}}\varepsilon_{0}.$
We arrive at
(31) $\operatorname{dist}_{a}\left(M_{h}(\lambda_{i}),U_{0}^{(i)}\right)\leqslant\sqrt{d_{i}}\varepsilon_{0}\leqslant\sqrt{D}\varepsilon_{0},\quad i=1,\ldots,q.$
For $(1,1)\leqslant(i,j)\leqslant(q,d_{q})$, we have from Eq. 29 that
$\left|\lambda^{h}_{ij}-\bar{\lambda}_{i}\right|\leqslant\left|\lambda^{h}_{ij}-\mathcal{C}_{i}\left(\left\{\lambda^{h}_{ij}\right\}_{j=1}^{d_{i}}\right)\right|+\left|\mathcal{C}_{i}\left(\left\{\lambda^{h}_{ij}\right\}_{j=1}^{d_{i}}\right)-\bar{\lambda}_{i}\right|\leqslant\gamma+\zeta_{0}<\frac{g}{2}.$
Consider $U_{1/2}^{(i)}=\operatorname{span}\{u_{i1}^{(1/2)},\ldots,u_{id_{i}}^{(1/2)}\}$ for $i=1,\ldots,q$. In accordance with Theorem 4.5 and Eq. 31, there holds that
$\operatorname{dist}_{a}\left(M_{h}(\lambda_{i}),U_{1/2}^{(i)}\right)\leqslant\frac{\sqrt{D}\left(\gamma+\zeta_{0}\right)\varepsilon_{0}}{\sqrt{(g-\gamma-\zeta_{0})^{2}(1-D\varepsilon_{0}^{2})+D\left(\gamma+\zeta_{0}\right)^{2}\varepsilon^{2}_{0}}},\quad i=1,\ldots,q.$
Since $\varepsilon_{0}$ is sufficiently small, we obtain from Proposition 3.5 that $U_{1}=\bigoplus_{i=1}^{q}U_{1/2}^{(i)}$, which together with Proposition 4.3 implies that
$\operatorname{dim}(U_{1})=\sum_{i=1}^{q}\operatorname{dim}(U_{1/2}^{(i)})=\sum_{i=1}^{q}d_{i}=N.$
Due to Eq. 18, we have
$\operatorname{dist}_{a}\left(\bigoplus_{i=1}^{q}M_{h}(\lambda_{i}),U_{1}\right)\leqslant\frac{\sqrt{DN}\left(\gamma+\zeta_{0}\right)\varepsilon_{0}}{\sqrt{(g-\gamma-\zeta_{0})^{2}(1-D\varepsilon_{0}^{2})+D\left(\gamma+\zeta_{0}\right)^{2}\varepsilon^{2}_{0}}}:\triangleq\xi_{1}.$
We apply Proposition 2.3 and Theorem 3.7 to Eq. 27 and obtain that there exists an orthonormal basis $\{u^{h,o,1}_{ij}\}_{j=1}^{d_{i}}$ of $M_{h}(\lambda_{i})$ with $b(u^{h,o,1}_{ij},u^{h,o,1}_{kl})=\delta_{ik}\delta_{jl}$ such that
(32) $\operatorname{dist}_{a}\left(u^{h,o,1}_{ij},u_{ij}^{(1)}\right)\leqslant C\xi_{1},\quad\lambda_{ij}^{(1)}-\lambda^{h}_{ij}\leqslant\lambda^{h}_{q+1,1}\xi_{1}^{2},\quad(1,1)\leqslant(i,j)\leqslant(q,d_{q}),$
where $C$ is a constant that is independent of $U_{1}$, i.e. independent of the iteration.
Set $\varepsilon_{1}=C\xi_{1}$ and $\zeta_{1}=\lambda^{h}_{q+1,1}\xi_{1}^{2}$, we then have
$\left|\mathcal{C}_{i}\left(\left\{\lambda^{h}_{ij}\right\}_{j=1}^{d_{i}}\right)-\bar{\lambda}_{i}^{(1)}\right|=\left|\mathcal{C}_{i}\left(\left\{\lambda^{h}_{ij}-\lambda_{ij}^{(1)}\right\}_{j=1}^{d_{i}}\right)\right|\leqslant\zeta_{1}.$
We obtain that $\varepsilon_{1}\leqslant\varepsilon_{0}$ and $\zeta_{1}\leqslant\zeta_{0}$ since $\varepsilon_{0},\gamma$ and $\zeta_{0}$ are sufficiently small. Similarly, there hold that
$\operatorname{dist}_{a}\left(\bigoplus_{i=1}^{q}M_{h}(\lambda_{i}),U_{n+1}\right)\leqslant$ $\xi_{n+1}=\frac{\sqrt{DN}\left(\gamma+\zeta_{n}\right)\varepsilon_{n}}{\sqrt{(g-\gamma-\zeta_{n})^{2}(1-D\varepsilon_{n}^{2})+D\left(\gamma+\zeta_{n}\right)^{2}\varepsilon^{2}_{n}}},$
$0\leqslant\lambda_{ij}^{(n+1)}-\lambda^{h}_{ij}\leqslant$ $\zeta_{n+1}=\lambda^{h}_{q+1,1}\xi_{n+1}^{2},\quad(1,1)\leqslant(i,j)\leqslant(q,d_{q}).$
Therefore there exists an orthonormal basis $\{u^{h,o,n+1}_{ij}\}_{j=1}^{d_{i}}$ of $M_{h}(\lambda_{i})$ such that
$\operatorname{dist}_{a}\left(u^{h,o,n+1}_{ij},u_{ij}^{(n+1)}\right)\leqslant\varepsilon_{n+1}=$ $C\xi_{n+1}=\frac{C\sqrt{DN}\left(\gamma+\zeta_{n}\right)\varepsilon_{n}}{\sqrt{(g-\gamma-\zeta_{n})^{2}(1-D\varepsilon_{n}^{2})+D\left(\gamma+\zeta_{n}\right)^{2}\varepsilon^{2}_{n}}},$
for $(1,1)\leqslant(i,j)\leqslant(q,d_{q})$, where the constant $C$ is the same as the one in Eq. 32 due to the independence of iterations.
We see that $\varepsilon_{n+1}\leqslant\varepsilon_{n}$ and $\zeta_{n+1}\leqslant\zeta_{n}$, and both $\varepsilon_{n}$ and $\zeta_{n}$ decrease towards 00 as $n\rightarrow\infty$. Finally, we arrive at
$\lim_{n\rightarrow\infty}\frac{\varepsilon_{n+1}}{\varepsilon_{n}}=\frac{C\sqrt{DN}\gamma}{g-\gamma}.$
Note that $\gamma\ll 1$ when $\operatorname{dist}_{a}\left(\bigoplus_{i=1}^{q}M(\lambda_{i}),V^{h}\right)\ll 1$, which together with Eq. 30 implies that Algorithm 3 converges faster when the finite dimensional discretization Eq. 3 approximates Eq. 1 better.
If Eq. 1 is already a discrete eigenvalue problem, then $\gamma=0$. We obtain from the proof of Theorem 4.7 that
(33) $\lim_{n\rightarrow\infty}\frac{\varepsilon_{n+1}}{\varepsilon_{n}^{3}}=\frac{\sqrt{DN}\lambda^{h}_{q+1,1}}{gC},$
which implies that it is a cubic convergence result. Note that above cubic convergence actually stems from the convergence of the shift parameters to some eigenvalues of (2.3), which is exactly what $\gamma=0$ means. The constant $C$ in Eq. 33 comes from the application of Theorem 3.7. Indeed, we can choose a larger $C$ since Eq. 33 implies that Algorithm 3 converges faster with larger $C$. However, when $C$ is larger, more exact initial values are required. For example, if we expect $\varepsilon_{n}$ to decrease towards 00 as $n\rightarrow\infty$, a necessary condition
$\varepsilon_{0}\geqslant\varepsilon_{1}=\frac{C\sqrt{DN}\zeta_{0}\varepsilon_{0}}{\sqrt{(g-\zeta_{0})^{2}(1-D\varepsilon_{0}^{2})+D\zeta_{0}^{2}\varepsilon^{2}_{0}}},$
implies that $\varepsilon_{0}$ and $\zeta_{0}$ are required smaller with larger $C$. Hence, from Theorem 3.7 and Eq. 30, we obtain that Algorithm 3 applied to a discrete eigenvalue problem converges in cubic rate and goes faster with more exact initial values. The classical result in the $1$-D discrete case $(N=1)$ stated as
$\lim_{n\rightarrow\infty}\frac{\varepsilon_{n+1}}{\varepsilon_{n}^{3}}\leqslant 1,$
under the assumption of convergence of the algorithm (see, e.g. [2, 23]). In contrast, Theorem 4.7 is more precise and also for general clustered eigenvalue problems. In addition, Eq. 30 and Eq. 33 show what the speed of the convergence depends on.
To improve the numerical stability, a modified shifted-inverse based ParO algorithm is proposed in [22]. We note that step $2$ of Algorithm 3 can also be written as follows: for $(1,1)\leqslant(i,j)\leqslant(q,d_{q})$, find $e_{ij}^{(n+1/2)}\in V^{h}$ in parallel satisfying
(34) $a(e_{ij}^{(n+1/2)},v)-\bar{\lambda}_{i}^{(n)}b(e_{ij}^{(n+1/2)},v)=2\bar{\lambda}_{i}^{(n)}b(u_{ij}^{(n)},v)-a(u_{ij}^{(n)},v)\quad\forall v\in V^{h},$
and set $u_{ij}^{(n+1/2)}=u_{ij}^{(n)}+e_{ij}^{(n+1/2)}$. Then instead of solving the $N$-dimensional projected eigenvalue problem in $U_{n+1}=\operatorname{span}\left\{u_{11}^{(n+1/2)},\ldots,u_{1d_{1}}^{(n+1/2)},\ldots,u_{qd_{q}}^{(n+1/2)}\right\}$, we consider the augmented $2N$-dimensional subspace
$\tilde{U}_{n+1}=\operatorname{span}\left\{u_{11}^{(n+1/2)},\ldots,u_{1d_{1}}^{(n+1/2)},\ldots,u_{qd_{q}}^{(n+1/2)},e_{11}^{(n+1/2)},\ldots,e_{1d_{1}}^{(n+1/2)},\ldots,e_{qd_{q}}^{(n+1/2)}\right\}.$
For the completeness of this paper, we show the modified shifted-inverse based ParO algorithm proposed in [22] here, which is stated as Algorithm 4.
Algorithm 4 Modified Shifted-Inverse Based ParO Algorithm
  1. 1.
    Given a finite-dimensional space $V^{h}$ and $tol>0$, provide and cluster initial data by Eq. 12, i.e., $\left\{\left(\lambda_{ij}^{(0)},u_{ij}^{(0)}\right)\right\}_{(1,1)\leqslant(i,j)\leqslant(q,d_{q})}\subset\mathbb{R}\times V^{h}$ with $b\left(u_{ij}^{(0)},u_{kl}^{(0)}\right)=\delta_{ik}\delta_{jl}$. Set $\bar{\lambda}_{i}^{(0)}=\mathcal{C}_{i}\left(\left\{\lambda_{ij}^{(0)}\right\}_{j=1}^{d_{i}}\right)$ and let $n=0$.
  2. 2.
    For $(1,1)\leqslant(i,j)\leqslant(q,d_{q})$, find $e_{ij}^{(n+1/2)}\in V^{h}$ in parallel by solving
    $a(e_{ij}^{(n+1/2)},v)-\bar{\lambda}_{i}^{(n)}b(e_{ij}^{(n+1/2)},v)=2\bar{\lambda}_{i}^{(n)}b(u_{ij}^{(n)},v)-a(u_{ij}^{(n)},v)\quad\forall v\in V^{h}.$
  3. 3.
    Construct $\tilde{U}_{n+1}=\operatorname{span}\left\{u_{11}^{(n+1/2)},\ldots,u_{qd_{q}}^{(n+1/2)},e_{11}^{(n+1/2)},\ldots,e_{qd_{q}}^{(n+1/2)}\right\}.$
  4. 4.
    Find $(\lambda^{(n+1)},u^{(n+1)})\in\mathbb{R}\times\tilde{U}_{n+1}$ satisfying
    (35) $a\left(u^{(n+1)},v\right)=\lambda^{(n+1)}b\left(u^{(n+1)},v\right)\quad\forall v\in\tilde{U}_{n+1},$
    to obtain eigenpairs $\left\{\left(\lambda_{ij}^{(n+1)},u_{ij}^{(n+1)}\right)\right\}$ with $b\left(u_{ij}^{(n+1)},u_{kl}^{(n+1)}\right)=\delta_{ik}\delta_{jl}$.
  5. 5.
    If $\sum_{i=1}^{q}\sum_{j=1}^{d_{i}}\left|\lambda_{ij}^{(n+1)}-\lambda_{ij}^{(n)}\right|>tol$, set $\bar{\lambda}_{i}^{(n+1)}=\mathcal{C}_{i}\left(\left\{\lambda_{ij}^{(n+1)}\right\}_{j=1}^{d_{i}}\right)$, $n=n+1$ and go to $2$.
The convergence of Algorithm 4 follows from the similar argument of the proof for Theorem 4.7 together with the fact that
$\tilde{U}_{n+1}=U_{n+1}\cup U_{n},\quad\operatorname{dist}_{a}(M_{h}(\lambda_{i}),\tilde{U}_{n+1})\leqslant\operatorname{dist}_{a}(M_{h}(\lambda_{i}),U_{n+1}).$
5 Concluding Remarks
In this paper, we have provided the numerical analysis of the parallel orbital-updating approach for linear eigenvalue problems based on the investigation of a quasi-orthogonality. Under the framework of the ParO approach, we have shown the convergence of some practical algorithms. We point out that numerical experiments in [9, 11, 22] show that the ParO approach is very efficient for electronic structure calculations. Due to the space limitation, we shall address the numerical analysis of the approach for the Kohn-Sham equation in a separate article. It is also our ongoing work to carry out the numerical analysis for the ParO based optimization approach proposed in [11].